### v2

@@ Appendix C Calculation of synchrotron radiation
where $e$ represents the total energy density in Eq. 3. For self-consistency, we adopt $\eta_{e}=0.1$ as we expect the non-thermal electrons to have negligible impact on the dynamics of the jets and lobes in our simulations. Finally, $N_{0}$ can be derived from Eqs. 35 and 36.
@@
where $l_{-}-l_{+}$ represents the integral path traversing the radio sources along the line of sight perpendicular to the jet axis. $\Delta A$ denotes the cross-sectional area of the simulation cells along the line of sight, and $d_{L}$ is the luminosity distance. For this study, we place the sources in our simulations at a redshift of $z=0.3$ (typical value of the sources in Dabhade et al. (2020b)) and calculate the radio power and flux at frequency $\nu=144\rm(1+z)MHz$. The radiating lobe region is identified by a density lower than $5\times 10^{-27}\rm{g\ cm^{-3}}$ and a toroidal magnetic field stronger than $10^{-9}\rm{G}$. In Figs. 4 and 9, we present snapshots of the radio flux distribution of the sources from selected runs, alongside the corresponding distributions of density and magnetic field.
@@
For a typical magnetic field of $B\sim 10^{-6}$ G and a frequency $\rm\nu=144(1+0.3)MHz$, Eq. 39 results in a cooling time about 1.4 Gyr, which is much longer than the simulation time for the most of our runs (500Myr) in this work.
@@
Appendix D Additional information on energy transport processes
Here, we present additional information, including the 2D plots that depict the total energy outflow flux at the surface of the jetted bubble for run t200fm8 (Fig. 10), the power of energy outflows (Fig. 11), and the volume-averaged energy density (Fig. 12) of various energy components varying with the traveling distance of the lobes in the two sets of runs discussed in Sects. 4.1 and 4.2.
Figure 10: Snapshots of the total energy outflow flux at the surface of the jetted bubble in run t200fm8 for different times.
Figure 11: Power of energy outflows for different energy components varying with the traveling distance of lobes in the two sets of runs discussed in Sect. 4.1. Different colors represent different runs, while the dotted, dashed, and solid lines denote kinetic, thermal, and magnetic energy outflows, respectively.

### v3

@@ Appendix C Calculation of synchrotron radiation
where $e$ represents the total energy density in Eq. 3. For self-consistency, we adopt $\eta_{e}=0.1$ as we expect the non-thermal electrons to have negligible impact on the dynamics of the jets and lobes in our simulations. Finally, $N_{0}$ can be derived from Eqs. 38 and 39. In Fig. 9, we show the volume-averaged energy density of various energy components varying with the traveling distance of the lobes.
@@
where $l_{-}-l_{+}$ represents the integral path traversing the radio sources along the line of sight perpendicular to the jet axis. $\Delta A$ denotes the cross-sectional area of the simulation cells along the line of sight, and $d_{L}$ is the luminosity distance. For this study, we place the sources in our simulations at a redshift of $z=0.3$ (typical value of the sources in Dabhade et al. (2020b)) and calculate the radio power and flux at frequency $\nu=144\rm(1+z)MHz$. The radiating lobe region is identified by a density lower than $5\times 10^{-27}\rm{g\ cm^{-3}}$ and a toroidal magnetic field stronger than $10^{-9}\rm{G}$. In Figs. 4 and 10, we present snapshots of the radio flux distribution of the sources from selected runs, alongside the corresponding distributions of density and magnetic field.
@@
For the frequency of $\rm\nu=144(1+0.3)MHz$, Eq. 42 results in a cooling time of about 43.9 Myr for a magnetic field of $10^{-5}$ G and 1.4 Gyr for $10^{-6}$ G, which means the electrons will undergo significant synchrotron cooling if the magnetic field is sufficiently strong when the jets are shut off.
@@
Appendix D Integral energy equation in Lagrangian form
Here we consider the energy loss from a given jetted bubble as it interacts with the surrounding medium. The energy equation Eq.3 can be written in its full conservation form, as usually shown in textbooks (Ogilvie 2016). Omitting the jet injection, we have:
$\frac{\partial}{\partial t}(e+\rho\Phi)+\nabla\cdot[(e+\rho\Phi)\textbf{v}+(p+\frac{B^{2}}{2})\textbf{v}-\textbf{B}\textbf{B}\cdot\textbf{v}]=0.$ (43)
When considering the energy outflow from a moving and growing bubble, we must refer to the integral energy equation in Lagrangian form. While neglecting the mixing processes, the energy outflow rate from the region $V_{bub}$ of the given bubble can be written as
[…]
Here we have used the relation between the rate of change of a small volume and velocity $\frac{1}{\Delta V}\frac{d\Delta V}{dt}=\nabla\cdot\textbf{v}$ and the relation between Lagrangian and Eulerian derivatives $\frac{d}{dt}=\frac{\partial}{\partial t}+\textbf{v}\cdot\nabla$ (or applying the Reynolds Transport Theorem). Then, using Eq.43, we have
[…]
where $S_{bub}$ is the surface of the bubble. The physical meaning is obvious: the bubble mainly loses energy through adiabatic processes due to the work done by the thermal pressure and Maxwell stress tensor (magnetic pressure plus magnetic tension). The same derivation for the mass equation results in a zero mass loss rate, indicating the mass conservation in the moving bubbles when the mixing processes are neglected.